# Masked Language Modeling Becomes Conditional Density Estimation for Tabular Data Synthesis

### v1

@@ A.1 Proof of Theorem 1
[…]
Then, by the Cauchy-Schwarz inequality,
$\mathbb{E}_{\mathbb{P}_{2}}\left[\left(\tilde{F}_{j}({\bf x}_{j})-F^{*}_{j}({\bf x}_{j})\right)^{2}\right]=\mathbb{E}_{\mathbb{P}_{2}}\left[\left(\int_{-\infty}^{{\bf x}_{j}}\tilde{p}_{j}(x)-p^{*}_{j}(x)dx\right)^{2}\right]$ (10)
$=$ $\mathbb{E}_{\mathbb{P}_{2}}\left[\left(\int_{-\infty}^{-M}\tilde{p}_{j}(x)-p^{*}_{j}(x)dx+\int_{-M}^{{\bf x}_{j}}\tilde{p}_{j}(x)-p^{*}_{j}(x)dx\right)^{2}\right]$
$\leq$ $\mathbb{E}_{\mathbb{P}_{2}}\left[2\cdot\left(\int_{-\infty}^{-M}\tilde{p}_{j}(x)-p^{*}_{j}(x)dx\right)^{2}+2\cdot\left(\int_{-M}^{{\bf x}_{j}}\tilde{p}_{j}(x)-p^{*}_{j}(x)dx\right)^{2}\right]$
[…]
By the sub-additivity,
$\mathbb{P}\Bigg{(}\Big{|}\hat{p}_{j}({\bf x}_{j}|{\bf x}_{-j};\theta)-p_{j}^{*}({\bf x}_{j}|{\bf x}_{-j})\Big{|}>\epsilon\Bigg{)}$
$=$ $\mathbb{P}\Bigg{(}\Big{|}\hat{c}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j};\theta)\cdot\tilde{p}_{j}({\bf x}_{j})-c^{*}_{j}(F^{*}_{j}({\bf x}_{j})|{\bf x}_{-j})\cdot p^{*}_{j}({\bf x}_{j})\Big{|}>\epsilon\Bigg{)}$
$\leq$ $\mathbb{P}\Bigg{(}\Big{|}\hat{c}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j};\theta)\cdot\tilde{p}_{j}({\bf x}_{j})-c^{*}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j})\cdot\tilde{p}_{j}({\bf x}_{j})\Big{|}>\frac{\epsilon}{2}\Bigg{)}\text{\quad$\cdots(A)$}$
$+\mathbb{P}\Bigg{(}\Big{|}c^{*}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j};\theta)\cdot\tilde{p}_{j}({\bf x}_{j})-c^{*}_{j}(F^{*}_{j}({\bf x}_{j})|{\bf x}_{-j})\cdot p^{*}_{j}({\bf x}_{j})\Big{|}>\frac{\epsilon}{2}\Bigg{)},\text{\quad$\cdots(B)$}$
where $\mathbb{P}$ denotes the probability range over different choices of all $n$ training set.
$(A)$ For any $\eta\in(0,1)$, there exists a sufficiently large $M^{*}>0$ such that
$\Pr\left(\{x:p_{j}^{*}(x)>M^{*}\}\right)<\eta.$
By the sub-additivity,
$\mathbb{P}\Bigg{(}\Big{|}\hat{c}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j};\theta)\cdot\tilde{p}_{j}({\bf x}_{j})-c_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j})\cdot\tilde{p}_{j}({\bf x}_{j})\Big{|}>\frac{\epsilon}{2}\Bigg{)}$
$=$ $\mathbb{P}\Bigg{(}\Big{|}\hat{c}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j};\theta)-c^{*}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j})\Big{|}\cdot\Big{|}\tilde{p}_{j}({\bf x}_{j})\Big{|}>\frac{\epsilon}{2}\Bigg{)}$
$\leq$ $\mathbb{P}\Bigg{(}\Big{|}\hat{c}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j};\theta)-c^{*}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j})\Big{|}\cdot\Big{|}\tilde{p}_{j}({\bf x}_{j})-p^{*}_{j}({\bf x}_{j})\Big{|}>\frac{\epsilon}{4}\Bigg{)}$
$+\mathbb{P}\Bigg{(}\Big{|}\hat{c}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j};\theta)-c^{*}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j})\Big{|}\cdot\Big{|}p^{*}_{j}({\bf x}_{j})\Big{|}>\frac{\epsilon}{4}\Bigg{)}$
$\leq$ $\mathbb{P}\Bigg{(}\Big{|}\hat{c}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j};\theta)-c^{*}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j})\Big{|}>\sqrt{\frac{\epsilon}{4}}\Bigg{)}+\mathbb{P}\Bigg{(}\Big{|}\tilde{p}_{j}({\bf x}_{j})-p^{*}_{j}({\bf x}_{j})\Big{|}>\sqrt{\frac{\epsilon}{4}}\Bigg{)}$
$+\mathbb{P}\Bigg{(}\Big{|}\hat{c}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j};\theta)-c^{*}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j})\Big{|}\cdot\Big{|}p^{*}_{j}({\bf x}_{j})\Big{|}>\frac{\epsilon}{4},\Big{\{}{\bf x}_{j}:p_{j}^{*}({\bf x}_{j})\leq M^{*}\Big{\}}\Bigg{)}$
$+\mathbb{P}\Bigg{(}\Big{|}\hat{c}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j};\theta)-c^{*}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j})\Big{|}\cdot\Big{|}p^{*}_{j}({\bf x}_{j})\Big{|}>\frac{\epsilon}{4},\Big{\{}{\bf x}_{j}:p_{j}^{*}({\bf x}_{j})>M^{*}\Big{\}}\Bigg{)}$
$\leq$ $\mathbb{P}\Bigg{(}\Big{|}\hat{c}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j};\theta)-c^{*}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j})\Big{|}>\sqrt{\frac{\epsilon}{4}}\Bigg{)}+\mathbb{P}\Bigg{(}\Big{|}\tilde{p}_{j}({\bf x}_{j})-p^{*}_{j}({\bf x}_{j})\Big{|}>\sqrt{\frac{\epsilon}{4}}\Bigg{)}$
$+\mathbb{P}\Bigg{(}\Big{|}\hat{c}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j};\theta)-c^{*}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j})\Big{|}\cdot M^{*}>\frac{\epsilon}{4}\Bigg{)}+\mathbb{P}\Big{(}\Big{\{}{\bf x}_{j}:p_{j}^{*}({\bf x}_{j})>M^{*}\Big{\}}\Big{)}.$
Thus, for every $\epsilon>0$,
$\lim_{n\rightarrow\infty}\mathbb{P}\Bigg{(}\Big{|}\hat{c}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j};\theta)\cdot\tilde{p}_{j}({\bf x}_{j})-c^{*}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j})\cdot\tilde{p}_{j}({\bf x}_{j})\Big{|}>\frac{\epsilon}{2}\Bigg{)}=0,$
because $\hat{c}_{j}$ is MSE-consistent estimator and $\tilde{p}_{j}$ is a kernel density estimator.
$(B)$ Furthermore, by the continuous mapping theorem and the algebra of the convergence in probability, for every $\epsilon>0$,
$\lim_{n\rightarrow\infty}\mathbb{P}\Bigg{(}\Big{|}c^{*}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j};\theta)\cdot\tilde{p}_{j}({\bf x}_{j})-c^{*}_{j}(F^{*}_{j}({\bf x}_{j})|{\bf x}_{-j})\cdot p^{*}_{j}({\bf x}_{j})\Big{|}>\frac{\epsilon}{2}\Bigg{)}=0$
because both $\tilde{F}_{j}$ and $\tilde{p}_{j}$ are MSE-consistent estimators.
Therefore, for every $\epsilon>0$,
$\lim_{n\rightarrow\infty}\mathbb{P}\Bigg{(}\Big{|}\hat{p}_{j}({\bf x}_{j}|{\bf x}_{-j};\theta)-p^{*}_{j}({\bf x}_{j}|{\bf x}_{-j})\Big{|}>\epsilon\Bigg{)}=0$
as $n\rightarrow\infty$. It means that $\hat{p}_{j}({\bf x}_{j}|{\bf x}_{-j};\theta)$ is weakly consistent estimator for $p^{*}_{j}({\bf x}_{j}|{\bf x}_{-j})$. The proof is complete. ∎
A.2 Proof of Proposition 1
@@
A.3 Dataset Descriptions
@@
A.4 Missing Mechanism
A.5 Experimental Settings for Reproduction
@@
  • In practice, for all $j\in I_{C}$, we estimate $\hat{F}_{j}$ using empirical measure. In the presence of missing data, $\hat{F}_{j}$ is estimated solely using the observed dataset (refer to [6]).
@@
  • Optimizer: AdamW optimizer [32] with 0.001 weight decay parameter.
  • The hyper-parameter search space for fitting MaCoDE is shown in Table 5.
Table 5: The hyper-parameter search space for MaCoDE. We utilized the best-performing combination of hyper-parameters within this search space. The chosen hyper-parameter is bolded.
hyper-parameter search space
epochs $\{100,200,\mathbf{500}\}$
batch size $\{512,\mathbf{1024}\}$
learning rate $\{0.0001,0.0005,\mathbf{0.001}\}$
$L$$\{30,\mathbf{50},100\}$
Transformer encoder dimension $\{\mathbf{128},256\}$
Transformer encoder #heads $\{2,\mathbf{4},8\}$
Transformer encoder #layer $\{1,\mathbf{2},4\}$
A.5.1 Details of Implementing Baseline Models
@@
  • CTGAN and TVAE [51]: We rely on the implementations provided in the official code  [^3] https://github.com/sdv-dev/CTGAN for CTGAN and TVAE, where hyperparameters are predefined. We adopt the default hyperparameters specified in the module. To ensure fairness in comparison, we adjust the latent dimension of CTGAN’s generator to 100. Similarly, we increase the sizes of the TVAE’s latent dimension to 100.
@@
  • CTAB-GAN and CTAB-GAN+ [56, 57]: We adhere to the official implementations [^4] https://github.com/Team-TUD/CTAB-GAN, https://github.com/Team-TUD/CTAB-GAN-Plus. The latent dimension is fixed at 100, and the maximum number of clusters is set to 10. Since the specification of mixed-type and general-type features is discretionary, we use continuous, integer, and categorical variables.
@@
  • DistVAE [1]: We follow the official implementations [^5] https://github.com/an-seunghwan/DistVAE. However, we changed the latent dimension to 100 for a fair comparison with other models.
@@
  • TabDDPM [24]: We utilized the TabDDPM module in synthcity.Plugins  [^6] https://github.com/vanderschaarlab/synthcity for synthetic data generation.
@@
  • TabMT [13]: Rather than using K-means clustering, we utilize the Gaussian Mixture Model (GMM) to discretize continuous columns with the objective of preserving the original continuous domain. We determine the optimal number of clusters for the GMM, ranging from 2 to 10, based on the Bayesian Information Criterion (BIC). During the generation of continuous columns, we initially predict the component label and then sample from the selected Gaussian component.
@@
  • MICE [48]: We employed the IterativeImputer package from Scikit-learn for multiple imputation using chained equations. Following the authors’ experiments, a max_iter range of 10 to 20 was considered sufficient for convergence, and we adopted this setting for our experiments. Additionally, to introduce randomness, we set imputation_order to random, which randomly selects a variable for imputation in each iteration. The remaining parameters were left at their default values to maintain the integrity of the MICE implementation.
@@
  • GAIN [52]: We adhere to the official implementations  [^7] https://github.com/jsyoon0823/GAIN/tree/master. As the paper does not explicitly discuss the separate handling of categorical and continuous variables, the code treats them simultaneously. Consequently, a rounding process is employed to handle categorical variables afterward.
  • missMDA [21]: We utilized the missMDA package [^8] https://cran.r-roject.org/web/packages/missMDA/index.html in R for multiple imputation. However, we encountered out-of-memory issues with covtype and letter datasets among the 10 datasets used in our experiments. As a result, we only present results for the remaining 8 datasets when evaluating the performance of missMDA. For details on the results related to Q3 using these 8 datasets, please refer to Table 18.
@@
  • VAEAC [19]: We adhere to the official implementations [^9] https://github.com/tigvarts/vaeac. The authors provided hyperparameters that adequately address both continuous and categorical variables, so we used these without further modification during model fitting.
@@
  • MIWAE [34]: The implemented MIWAE code [^10] https://github.com/pamattei/miwae was designed for continuous variables only. To accommodate heterogeneous tabular datasets, we treated the conditional distribution of categorical columns as categorical distributions and employed cross-entropy loss for reconstruction. For comparison with not-MIWAE, we set the latent dimension to $p-1$.
@@
  • not-MIWAE [18]: The implemented not-MIWAE code [^11] https://github.com/nbip/notMIWAE also focused on continuous variables exclusively. To handle categorical variables, we made the same modifications as in MIWAE. For comparison with MIWAE, we set the latent dimension to $p-1$. Training was conducted for 100K steps, consistent with the official implementations.
  • EGC [55]: We utilized the gcimpute package [^12] https://github.com/udellgroup/gcimpute/tree/master to implement EGC. However, we encountered difficulties in fitting EGC to concrete, kings, and loan datasets among the 10 datasets used in our experiments. Therefore, for the evaluation of EGC performance, we only provide results for the remaining 7 datasets. Please refer to Table 19 for the outcomes related to Q3 using these 7 datasets.
Table 6: The number of model parameters.
Model abalone banknote breast concrete covertype kings letter loan redwine whitewine
CTGAN 0.7M 0.4M 0.8M 0.5M 0.7M 1.1M 1.1M 0.5M 0.6M 0.7M
TVAE 0.2M 0.2M 0.2M 0.2M 0.2M 0.3M 0.3M 0.2M 0.2M 0.2M
CTAB-GAN 1.7M 0.5M 1.7M 1.1M 1.8M 1.8M 1.9M 1.1M 1.4M 1.7M
CTAB-GAN+ 1.7M 0.5M 1.8M 1.1M 1.8M 1.9M 1.9M 1.1M 1.1M 1.7M
DistVAE 0.2M 0.2M 0.3M 0.2M 0.2M 0.2M 0.2M 0.2M 0.2M 0.2M
TabDDPM 0.2M 0.2M 0.2M 0.2M 0.2M 0.2M 0.2M 0.2M 0.2M 0.2M
TabMT 1.2M 1.2M 1.2M 1.2M 1.2M 1.2M 1.2M 1.2M 1.3M 1.2M
GAIN 0.8K 0.2K 7.8K 0.7K 1.0K 2.7K 2.4K 1.0K 1.2K 1.2K
VAEAC 2.8M 2.8M 2.9M 2.8M 2.8M 2.9M 2.8M 2.8M 2.8M 2.8M
MIWAE 44.2K 37.3K 60.7K 42.4K 40.0K 55.8K 51.2K 42.3K 44.1K 44.2K
not-MIWAE 43.7K 36.8K 57.9K 41.6K 42.2K 55.9K 49.9K 41.9K 42.9K 43.0K
EGC 0.04K 0.01K 0.47K - 0.06K - 0.14K - 0.07K 0.07K
MaCoDE 1.3M 1.2M 1.6M 1.3M 1.3M 1.4M 1.4M 1.3M 1.3M 1.3M
A.5.2 Evaluation Settings for Q1, Q2 and Q3
  • Evaluation procedure for Q1: Regarding machine learning utility, we conduct a regression and classification task for data replacement evaluation.
    We assess regression prediction performance using the average SMAPE values of all continuous columns as the target variable once each with a Random Forest regressor. And we assess classification prediction performance for the classification target variable (see Table 4) using the average $F_{1}$ values of five different classifiers: Logistic Regression, Gaussian Naive Bayes, K-Nearest Neighbors classifier, Decision Tree classifier, and Random Forest classifier (refer to Table 7 for detailed configuration).
    For model selection evaluation, we train classifiers on both original training and synthetic datasets and evaluate their classification performance on the test dataset. We assess their effectiveness by comparing their AUROC rank orderings using Spearman’s Rank Correlation.
    Similarly, for feature selection evaluation, we train a random forest classifier on the original training dataset to determine the rank-ordering of important features. Then, we compare it with the rank ordering from the same model type trained on synthetic data using Spearman’s Rank Correlation.
  • Evaluation procedure for Q2: Our evaluation process is based on [13, 15]. The dataset is first split into training and testing sets across 10 different random seeds. For each seed, we generate the mask. Subsequently, the model is fitted with a masked training dataset and generates synthetic data. This synthetic data is then assessed based on its performance in a downstream task.
    For regression tasks, we evaluate the synthesizers using the average SMAPE values of all continuous columns as the target variable once each with a Random Forest regressor. For classification tasks, we evaluate the synthesizers the average $F_{1}$ values of five different classifiers for the classification target variable (see Table 4): Logistic Regression, Gaussian Naive Bayes, K-Nearest Neighbors classifier, Decision Tree classifier, and Random Forest classifier (refer to Table 7 for detailed configuration). The final step aggregates these metrics across all seeds, presenting an average along with error bars.
  • Evaluation procedure for Q3: We assess the effectiveness of multiple imputations by employing interval inference for the population mean, which was proposed by Rubin [41]. However, since obtaining the population mean is not feasible, we utilize the sample column-wise mean of the complete dataset as a parameter of interest [26, 53]. The evaluation procedure for multiple imputations is outlined in Algorithm 3, and we report the mean and standard error of bias, coverage, and confidence interval length across 10 different random seeds, all continuous columns, and datasets [47]. Note that we do not split the data into training and test sets [18].
    Following [36, 20, 54], we generate the missing value mask for each dataset with three mechanisms in four settings. (MCAR) In the MCAR setting, each value is masked according to the realization of a Bernoulli random variable with a fixed parameter. (MAR) In the MAR setting, for each experiment, a fixed subset of variables that cannot have missing values is sampled. Then, the remaining variables have missing values according to a logistic model with random weights, which takes the non-missing variables as inputs. A bias term is fitted using line search to attain the desired proportion of missing values. (MNAR) Finally, two different mechanisms are implemented in the MNAR setting. The first, MNARL, is identical to the previously described MAR mechanism, but the inputs of the logistic model are then masked by an MCAR mechanism. Hence, the logistic model’s outcome depends on missing values. The second mechanism, MNARQ, samples a subset of variables whose values in the lower and upper $p$th percentiles are masked according to a Bernoulli random variable, and the values in-between are left not missing.
@@ A.5.2 Evaluation Settings for Q1, Q2 and Q3
Algorithm 3 Evaluation procedure for multiple imputation [47, 41]

### v2

@@ Appendix A Appendix / supplemental material
$\leq$ $\frac{hK_{j}}{2}+\frac{1}{\sqrt{2}h}\left(\sum_{l=1}^{L}\pi_{jl}^{*}({\bf x}_{-j})\log\pi_{jl}^{*}({\bf x}_{-j})-\mathbb{E}_{{\bf y}_{j}|{\bf x}_{-j}}\left[\sum_{l=1}^{L}\mathbb{I}({\bf y}_{j}=l)\log\pi_{jl}(g({\bf x};F)\odot{\bf m}^{(j)};\theta)\right]\right)^{1/2}$
$=$ $\frac{hK_{j}}{2}+\frac{\sqrt{Bias(\theta)}}{\sqrt{2}h}$
$=$ $\frac{K_{j}}{2L}+\frac{\sqrt{Bias(\theta)}}{\sqrt{2}/L},$
where the inequality holds for (7).
The proof is complete. ∎
Proof of Proposition 2
@@
Dataset Descriptions
Download links.
  • abalone (Nash et al. 1995): https://archive.ics.uci.edu/dataset/1/abalone
  • banknote (Lohweg 2013): https://archive.ics.uci.edu/dataset/267/banknote+authentication
  • breast (Wolberg et al. 1995):
    https://archive.ics.uci.edu/dataset/17/breast+cancer+wisconsin+diagnostic
  • concrete (Yeh 2007):
    https://archive.ics.uci.edu/dataset/165/concrete+compressive+strength
  • covertype (Blackard 1998): https://www.kaggle.com/datasets/uciml/forest-cover-type-dataset
  • kings (CC0: Public Domain): https://www.kaggle.com/datasets/harlfoxem/housesalesprediction
  • letter (Slate 1991): https://archive.ics.uci.edu/dataset/59/letter+recognition
  • loan (CC0: Public Domain): https://www.kaggle.com/datasets/teertha/personal-loan-modeling
  • redwine (Cortez et al. 2009): https://archive.ics.uci.edu/dataset/186/wine+quality
  • whitewine (Cortez et al. 2009): https://archive.ics.uci.edu/dataset/186/wine+quality
@@
Experimental Settings for Reproduction
@@
  • In practice, for all $j\in I_{C}$, we estimate $\hat{F}_{j}$ using empirical measure. In the presence of missing data, $\hat{F}_{j}$ is estimated solely using the observed dataset (refer to (Chenouri, Mojirsheibani, and Montazeri 2009)).
Hyper-parameters of MaCoDE: As shown in Section A, MaCoDE consistently generates high-quality synthetic data without requiring an extensive hyperparameter tuning process, unlike methods such as (Kotelnikov et al. 2023; Lee, Kim, and Park 2023; Kim, Lee, and Park 2023). This demonstrates the generalizability of our proposed model to various tabular datasets. Our implementation codes for the proposed model, MaCoDE, are provided in the supplementary material. For all tabular datasets, we applied the following hyperparameters uniformly without any additional tuning:
  • epochs: 500
  • batch size: 1024
@@
  • learning rate: 0.001 (with AdamW optimizer (Loshchilov and Hutter 2017) with 0.001 weight decay parameter)
  • the number of bins: $L=50$
  • Transformer encoder dimension: 128
  • Transformer encoder #heads: 4
  • Transformer encoder #layer: 2
  • Transformer encoder dropout ratio: 0.0
Details of Implementing Baseline Models
To compare our proposed method with baseline models, we performed experiments in a manner similar to (Zhang et al. 2024). We adjusted the hidden or latent dimensions across various methods, ensuring that the number of trainable parameters is comparable. Since each model employs distinct neural network architectures, it would be misleading to compare their performances using the same network configuration. It’s worth noting that model size typically reflects the performance of the baseline models. Under these conditions, we reproduced the baseline methods using their official codes and default configurations (except for the hidden or latent dimensions). Our reproduced codes for the baseline models are provided in the supplementary material. Below are the detailed implementations of the baseline methods:
@@
  • CTGAN and TVAE (Xu et al. 2019): We follow the implementations provided in the official repository [^4] https://github.com/sdv-dev/CTGAN for CTGAN and TVAE. We adopt the default hyperparameters specified in the module. To ensure fairness in comparison by aligning model sizes, we adjust the latent dimension of CTGAN and TVAE to 100.
@@
  • CTAB-GAN and CTAB-GAN+ (Zhao et al. 2021, 2023b): We follow the implementations provided in the official repository [^5] https://github.com/Team-TUD/CTAB-GAN, https://github.com/Team-TUD/CTAB-GAN-Plus. The latent dimension is set to 100, and the maximum number of clusters is set to 10. For the specification of feature types, we use three categories: continuous, integer, and categorical variables.
@@
  • DistVAE (An and Jeon 2023): We follow the implementations provided in the official repository [^6] https://github.com/an-seunghwan/DistVAE. We changed the latent dimension to 100 to increase the model size.
@@
  • TabDDPM (Kotelnikov et al. 2023): We utilized the TabDDPM module in synthcity.Plugins  [^7] https://github.com/vanderschaarlab/synthcity for synthetic data generation. This module follows the implementations provided in the official repository.
@@
  • TabMT (Gulati and Roysdon 2023): Rather than using K-means clustering, we utilize the Gaussian Mixture Model (GMM) to discretize continuous columns to preserve the original continuous domain. We determine the optimal number of clusters for the GMM, ranging from 2 to 10, based on the Bayesian Information Criterion (BIC). During the generation of continuous columns, we initially predict the component label and then sample from the selected Gaussian component.
@@
  • MICE (van Buuren and Groothuis-Oudshoorn 2011): We employed the IterativeImputer package from Scikit-learn for multiple imputation using chained equations. Following the authors’ experiments, a max_iter range of 10 to 20 was considered sufficient for convergence, and we adopted this setting for our experiments. Additionally, to introduce randomness, we set imputation_order to random, which randomly selects a variable for imputation in each iteration. The remaining parameters were left at their default values to maintain the integrity of the MICE implementation.
@@
  • GAIN (Yoon, Jordon, and van der Schaar 2018): We follow the implementations provided in the official repository [^8] https://github.com/jsyoon0823/GAIN/tree/master. As the paper does not explicitly discuss the separate handling of categorical and continuous variables, the code treats them simultaneously. Consequently, a rounding process is employed to handle categorical variables afterward.
  • missMDA (Josse and Husson 2016): We utilized the missMDA package [^9] https://cran.r-project.org/web/packages/missMDA/index.html in R for multiple imputation.
@@
  • VAEAC (Ivanov, Figurnov, and Vetrov 2019): We follow the implementations provided in the official repository [^10] https://github.com/tigvarts/vaeac. The authors provided hyperparameters that adequately address both continuous and categorical variables, so we used these without further modification during model fitting.
@@
  • MIWAE (Mattei and Frellsen 2019): The implemented MIWAE code in the official repository [^11] https://github.com/pamattei/miwae was designed for continuous variables only. To accommodate heterogeneous tabular datasets, we treated the conditional distribution of categorical columns as categorical distributions and employed cross-entropy loss for reconstruction. For comparison with not-MIWAE, we set the latent dimension to $p-1$.
@@
  • not-MIWAE (Ipsen, Mattei, and Frellsen 2021): The implemented not-MIWAE code in the official repository [^12] https://github.com/nbip/notMIWAE also focused on continuous variables exclusively. To handle categorical variables, we made the same modifications as in MIWAE. For comparison with MIWAE, we set the latent dimension to $p-1$. Training was conducted for 100K steps, consistent with the official implementations.
  • EGC (Zhao, Townsend, and Udell 2022): We utilized the gcimpute package [^13] https://github.com/udellgroup/gcimpute/tree/master to implement EGC. The model is free of hyperparameters.
@@
Evaluation Procedure of Q1
Regression performance (SMAPE).
  1. 1.
    Train a synthesizer using the real training dataset.
  2. 2.
    Generate a synthetic dataset with the same size as the real training dataset.
  3. 3.
    Train a machine learning model (Random Forest regressor) using the synthetic dataset, where each continuous column serves as the regression target variable.
  4. 4.
    Assess regression prediction performance by averaging the SMAPE values from the test dataset for each Random Forest regressor trained on the continuous columns.
Classification performance ($F_{1}$).
  1. 1.
    Train a synthesizer using the real training dataset.
  2. 2.
    Generate a synthetic dataset with the same size as the real training dataset.
  3. 3.
    Train machine learning models (Logistic Regression, Gaussian Naive Bayes, K-Nearest Neighbors classifier, Decision Tree classifier, and Random Forest classifier) using the synthetic dataset (see Table 3 for the classification target variable and refer to Table 4 for detailed configuration).
  4. 4.
    Assess classification prediction performance by averaging the $F_{1}$ values from the test dataset from five different classifiers.
Model selection performance (Model).
  1. 1.
    Train a synthesizer using the real training dataset.
  2. 2.
    Generate a synthetic dataset with the same size as the real training dataset.
  3. 3.
    Train machine learning models (Logistic Regression, Gaussian Naive Bayes, K-Nearest Neighbors classifier, Decision Tree classifier, and Random Forest classifier) using both the real training dataset and the synthetic dataset (see Table 3 for the classification target variable and refer to Table 4 for detailed configuration).
  4. 4.
    Evaluate the classification performance (AUROC) of all trained classifiers on the test dataset.
  5. 5.
    Assess model selection performance by comparing the AUROC rank orderings of classifiers trained on the real training dataset and those trained on the synthetic dataset using Spearman’s Rank Correlation.
Feature selection performance (Feature).
  1. 1.
    Train a synthesizer using the real training dataset.
  2. 2.
    Generate a synthetic dataset with the same size as the real training dataset.
  3. 3.
    Train a Random Forest classifier using both the real training dataset and the synthetic dataset (see Table 3 for the classification target variable and refer to Table 4 for detailed configuration).
  4. 4.
    Determine the rank-ordering of important features for both classifiers.
  5. 5.
    Assess feature selection performance by comparing the feature importance rank orderings of classifiers trained on the real training dataset and those trained on the synthetic dataset using Spearman’s Rank Correlation.
Evaluation Procedure of Q2
Regression performance (SMAPE).
  1. 1.
    For each random seed, we generate the mask and train MaCoDE using a masked training dataset (i.e., incomplete dataset).
  2. 2.
    Generate a synthetic dataset with the same size as the real training dataset.
  3. 3.
    Train a machine learning model (Random Forest regressor) using the synthetic dataset, where each continuous column serves as the regression target variable.
  4. 4.
    Assess regression prediction performance by averaging the SMAPE values from the test dataset for each Random Forest regressor trained on the continuous columns.
Classification performance ($F_{1}$).
  1. 1.
    For each random seed, we generate the mask and train MaCoDE using a masked training dataset (i.e., incomplete dataset).
  2. 2.
    Generate a synthetic dataset with the same size as the real training dataset.
  3. 3.
    Train machine learning models (Logistic Regression, Gaussian Naive Bayes, K-Nearest Neighbors classifier, Decision Tree classifier, and Random Forest classifier) using the synthetic dataset (see Table 3 for the classification target variable and refer to Table 4 for detailed configuration).
  4. 4.
    Assess classification prediction performance by averaging the $F_{1}$ values from the test dataset from five different classifiers.
Evaluation of Q3
Related Works
Recent methods employ the deep generative model to estimate a joint distribution and generate samples for imputations. GAN-based imputers such as GAIN (Yoon, Jordon, and van der Schaar 2018) and MisGAN (Li, Jiang, and Marlin 2019) adopt an adversarial learning approach to generate both missing entries and masking vectors. Other approaches like VAEAC (Ivanov, Figurnov, and Vetrov 2019), HI-VAE (Nazábal et al. 2020), and ReMasker (Du, Melis, and Wang 2024) employ strategies that allow them to learn conditional distributions on arbitrary conditioning sets using the uniform masking strategy. MIWAE (Mattei and Frellsen 2019), based on the Importance Weighted Autoencoder (Burda, Grosse, and Salakhutdinov 2015), demonstrates that under mild conditions and the MAR assumption, the target likelihood can be approximated regardless of the imputation function. Additionally, not-MIWAE (Ipsen, Mattei, and Frellsen 2021) extends MIWAE to handle cases where the data is under MNAR assumption by modeling missing entries as a latent variable. In parallel, the Optimal Transport (OT)-based method employs distributional matching, utilizing the 2-Wasserstein distance to compare distributions in both data and latent spaces (Muzellec et al. 2020; Zhao et al. 2023a). To handle mixed-type tabular datasets, (Zhao, Townsend, and Udell 2022) introduced EGC (extended Gaussian copula), which relies on a latent Gaussian distribution to support single and multiple imputations.
Evaluation Metrics
Algorithm 2 Evaluation procedure for multiple imputation (van Buuren 2012; Rubin and Schenker 1986)